# Image Can Bring Your Memory Back: A Novel Multi-Modal Guided Attack against Image Generation Model Unlearning

### v1

@@
Abstract.
Recent advances in image generation models (IGMs), particularly diffusion-based architectures such as Stable Diffusion (SD), have markedly enhanced the quality and diversity of AI-generated visual content. However, their generative capability has also raised significant ethical, legal, and societal concerns, including the potential to produce harmful, misleading, or copyright-infringing content. To mitigate these concerns, machine unlearning (MU) emerges as a promising solution by selectively removing undesirable concepts from pretrained models. Nevertheless, the robustness and effectiveness of existing unlearning techniques remain largely unexplored, particularly in the presence of multi-modal adversarial inputs.
To bridge this gap, we propose Recall, a novel adversarial framework explicitly designed to compromise the robustness of unlearned IGMs. Unlike existing approaches that predominantly rely on adversarial text prompts, Recall exploits the intrinsic multi-modal conditioning capabilities of diffusion models by efficiently optimizing adversarial image prompts with guidance from a single semantically relevant reference image. Extensive experiments across ten state-of-the-art unlearning methods and diverse tasks show that Recall consistently outperforms existing baselines in terms of adversarial effectiveness, computational efficiency, and semantic fidelity with the original textual prompt. These findings reveal critical vulnerabilities in current unlearning mechanisms and underscore the need for more robust solutions to ensure the safety and reliability of generative models. Code and data are publicly available at https://github.com/ryliu68/RECALL.
@@ Abstract.
Adversarial attacks, Image generation model unlearning, AI safety, Stable Diffusion model, AIGC
conference: ACM Conference; 2025journalyear: ;isbn:
1. Introduction
Figure 1. Given an unlearned IGM $\mathcal{G}_{u}$ that is assumed to have successfully eliminated the target content (e.g., nudity), our adversarial image prompt $P_{image}^{adv}$ combined with the original sensitive text prompt $P_{text}$ as multi-modal guidance, can still effectively circumvent the unlearning mechanism, leading to the reappearance of the removed content in the generated image $I^{*}$. Sensitive parts are covered by Refer to caption.
@@ 1. Introduction
The emergence of image generation models (IGMs), exemplified by Stable Diffusion (SD) (Rombach et al., 2022), has significantly enhanced the quality and variety of AI-generated visual content. These models have been successfully employed across various domains, including digital art creation, multimedia generation, and visual storytelling (Weng et al., 2024; Chen et al., 2024; Jiang et al., 2024; Zhang et al., 2024a), boosting the innovations for creative professionals. Nonetheless, their rapid advancement has simultaneously intensified ethical, societal, and legal concerns, specifically regarding potential misuse in generating harmful, misleading, or copyright-infringing content (Zhang et al., 2024b; Qu et al., 2023; Schramowski et al., 2023; Rando et al., 2022). Consequently, ensuring robust safety and trustworthiness mechanisms within these generative frameworks has emerged as an urgent imperative.
@@
Among different lines of efforts, machine unlearning (MU) has recently gained growing prominence (Gandikota et al., 2023; Zhang et al., 2024; Park et al., 2024; Li et al., 2024). It aims to remove sensitive concepts (e.g., nudity, violence, and copyrighted materials) from the IGMs, prohibiting the generation of sensitive or problematic content while maintaining the model’s general capability of producing benign and high-quality outputs (Schramowski et al., 2023; Kumari et al., 2023; Gandikota et al., 2024). Recent IGM unlearning (IGMU) methods utilize diverse strategies, including fine-tuning (Gandikota et al., 2023; Zhang et al., 2024b), targeted concept removal (Gandikota et al., 2024; Gong et al., 2024; Orgad et al., 2023), negative prompting (Schramowski et al., 2023), and adversarial filtering (Zhang et al., 2024; Gong et al., 2024; Wu et al., 2025). They have proven effective in safety protection of contemporary IGMs, enforcing compliance with ethical guidelines and legal standards.
@@
Despite the rapid progress in this field, the practical robustness of these techniques is challenged, especially under adversarial scenarios. Recent studies have revealed that unlearned IGMs are still vulnerable: carefully optimized prompts can successfully circumvent safety mechanisms, prompting the unlearned models to regenerate prohibited content (Chin et al., 2024; Zhang et al., 2024; Han et al., 2024; Tsai et al., 2024). However, these attack methods mainly focus on perturbing the textual modality and suffer from the following critical limitations. ① Modifying textual inputs can disrupt the semantic alignment between the generated images and original prompts; ② Many approaches rely on external classifiers or additional diffusion models for adversarial text prompt optimization, incurring substantial computational overhead; ③ Their effectiveness sharply declines against robust, adversarially-enhanced unlearning methods (e.g., AdvUnlearn (Zhang et al., 2024), RECE (Gong et al., 2024)); ④ Crucially, these methods overlook the inherent multi-modal conditioning capabilities (e.g., simultaneous textual and image) of IGMs, thus missing a critical dimension of potential vulnerability.
@@
To evaluate the vulnerability of existing unlearning methods and the effectiveness of our multi-modal attack Recall, we conduct extensive experiments involving ten state-of-the-art IGMU methods across four representative unlearning scenarios. Empirical results demonstrate that Recall consistently surpasses existing text-based adversarial prompting methods in terms of attack performance, computational efficiency, and semantic fidelity. These findings reveal critical vulnerabilities in current unlearning pipelines, highlighting their susceptibility to multi-modal guided adversarial attacks and underscoring the urgent need for developing more robust and verifiable unlearning mechanisms for image generation models. Our key contributions are as follows:
@@
  • We propose Recall, the first multi-modal guided adversarial attack framework to break the robustness of IGMU techniques, allowing the protected model to regenerate unlearned sensitive concepts with high semantic fidelity.
@@
  • Recall introduces a highly efficient optimization strategy that operates solely within the unlearned model by utilizing only a single reference image, eliminating the need for auxiliary classifiers, original diffusion models, or external semantic guidance (e.g., CLIP) required by previous attacks.
@@
  • Through comprehensive experiments covering ten representative IGMU techniques across four diverse unlearning tasks, we empirically demonstrate the vulnerabilities of existing unlearning solutions under multi-modal attacks, revealing the urgent need for more robust safety unlearning.
2. Related Work
2.1. Image Generation Model
@@
Image generation models (IGMs), particularly those based on diffusion architecture, have garnered substantial attention due to their ability to synthesize diverse, high-fidelity images via iterative denoising processes. Representative models include Stable Diffusion (SD) (Rombach et al., 2022), DALL·E (OpenAI, 2023), and Imagen (Saharia et al., 2022). These models typically leverage large-scale datasets (e.g., LAION-5B (Schuhmann et al., 2022)) and sophisticated architectural components, including: ① pre-trained text encoders (e.g., CLIP (Radford et al., 2021)), ② U-Net-based denoising backbones, and ③ VAE-based decoders. The integration of these components enables accurate semantic interpretation of textual inputs, facilitating controllable generation across diverse applications—from artistic expression to photorealistic synthesis.
2.2. Image Generation Model Unlearning
The widespread deployment of IGMs has also raised growing ethical and legal concerns, particularly their misuse in generating harmful, inappropriate, misleading, or copyrighted content (Qu et al., 2023; Li et al., 2024; Liu et al., 2025). To mitigate these risks, machine unlearning (MU) has been recently introduced as a lightweight manner to selectively remove specific concepts, styles, or objects from pretrained IGMs (Kumari et al., 2023; Wu et al., 2024; Fan et al., 2024), without affecting the overall generative capabilities. Existing image generation model unlearning (IGMU) techniques can be broadly categorized into three paradigms. ① Fine-tuning-based Unlearning: these methods adjust model parameters to erase specific learned representations. For example, Erased Stable Diffusion (ESD) (Gandikota et al., 2023) selectively fine-tunes the U-Net to suppress undesired features, while UCE (Gandikota et al., 2024) removes concepts via closed-form attention layer editing without further training. ② Guidance-based Unlearning: these approaches impose inference-time constraints without modifying model weights. Typical examples include negative prompt filtering and Safe Latent Diffusion (SLD) (Schramowski et al., 2023), which manipulate latent representations to block restricted content, offering high efficiency. ③ Regularization-based Knowledge Erasure: these strategies integrate forgetting signals into training objectives. For instance, Receler (Huang et al., 2024) applies contrastive regularization to suppress concept retention, and FMN (Zhang et al., 2024b) incorporates targeted regularizers to enforce structured forgetting during continued training.
2.3. Adversarial Attacks against IGMU
While IGMU techniques show promising results under standard conditions, their robustness against adversarial prompts remains largely uncertain. Recent studies have uncovered critical vulnerabilities, showing that carefully crafted prompts can bypass unlearning defenses and regenerate prohibited content (Li et al., 2024; Chin et al., 2024; Zhang et al., 2024). To explore these weaknesses, a variety of attack strategies have been proposed. Prompting4Debugging (P4D) (Chin et al., 2024) leverages CLIP and original SD models to optimize adversarial text prompts, achieving high attack success but incurring high computational cost. UnlearnDiffAtk (Zhang et al., 2024) improves efficiency by exploiting the internal discriminative capability of SD for direct prompt optimization. PUND (Han et al., 2024) constructs transferable embeddings via surrogate diffusion models for black-box attack settings. Ring-A-Bell (Tsai et al., 2024) identifies latent concept representations to generate prompts without model access. DiffZOO (Dang et al., 2024) applies zeroth-order optimization to perturb prompts in fully black-box settings. JPA (Ma et al., 2024) crafts discrete prefix prompts to bypass filters, while ICER (Chin et al., 2024) uses LLM-guided bandit optimization for interpretable black-box attacks.
Despite recent progress, existing attack methods face several key limitations. They often rely on text-based perturbations, which degrade semantic alignment between prompts and generated images. Many approaches also require external classifiers or access to the original diffusion model, increasing complexity and limiting scalability. Moreover, their effectiveness declines against robust unlearning techniques such as AdvUnlearn (Zhang et al., 2024), SafeGen (Li et al., 2024), and RECE (Gong et al., 2024). Finally, the adversarial optimization process is typically computationally intensive, limiting practical deployment.
@@ 2.3. Adversarial Attacks against IGMU
Therefore, an effective attack strategy must recover restricted content efficiently, maintain prompt-image semantic coherence, and exploit vulnerabilities beyond text perturbation. To this end, we propose Recall, a multi-modal adversarial framework that utilizes adversarial image prompts with unmodified text inputs to attack unlearned models using the multi-modal guidance. Our approach requires no external classifier or access to the original IGM, making it both lightweight and highly effective.
3. Preliminary
3.1. Stable Diffusion (SD) Model
SD models are grounded in denoising diffusion probabilistic models (DDPMs) (Ho et al., 2020), which operate via two complementary phases: a forward noising process and a reverse denoising process. The forward diffusion progressively corrupts the original data by iteratively adding Gaussian noise, formally characterized as:
(1) $q(z_{t}\mid z_{t-1})=\mathcal{N}(z_{t};\sqrt{\alpha_{t}}\,z_{t-1},(1-\alpha_{t})I),$
where $z_{t}$ denotes the latent representation at diffusion step $t$, and $\alpha_{t}$ dictates the variance schedule. After a fixed number of timesteps (typically 1,000), the input is gradually transformed into an isotropic Gaussian distribution.
Conversely, the reverse diffusion seeks to reconstruct the original data distribution by estimating the noise introduced at each diffusion step. This estimation is implemented through a neural network, parameterized by $\epsilon_{\theta}(z_{t},c,t)$, conditioned on auxiliary inputs $c$ (e.g., textual embeddings or image features), enabling multi-modal guidance. The associated training objective is defined as:
(2) $\mathcal{L}_{DM}=\mathbb{E}_{z_{t}\sim q,c,\epsilon\sim\mathcal{N}(0,I),t}\left[\|\epsilon-\epsilon_{\theta}(z_{t},c,t)\|_{2}^{2}\right].$
To improve efficiency, Latent Diffusion Models (LDMs) (Rombach et al., 2022) perform the diffusion process in a compressed latent space rather than directly in the pixel space, significantly reducing the computational cost. Building upon LDMs, SD (Rombach et al., 2022) further incorporates optimized components, including a robust CLIP-based text encoder (Radford et al., 2021) and larger, more diverse training datasets (Schuhmann et al., 2022), thereby achieving superior generation performance. Due to its balance of efficiency and generation quality, SD has become a widely adopted backbone in contemporary image synthesis tasks.
3.2. Image Generation Model Unlearning
@@
Given an image generation model (IGM) $\mathcal{G}$ trained over a rich concept space $\mathcal{C}$, Image Generation Model Unlearning (IGMU) aims to selectively eliminate the model’s ability to generate content associated with a specific set of sensitive concepts $\mathcal{C}^{\prime}\in\mathcal{C}$, while preserving its generative capabilities on the remaining concept space $\mathcal{C}\setminus\mathcal{C}^{\prime}$.
Formally, let $P_{text}$ denote a text prompt associated with a target concept $c\in\mathcal{C}^{\prime}$. The unlearning process is modeled by an algorithm $\mathcal{A}_{u}$ that modifies either the parameters or architecture of the generative model:
(3) $\mathcal{G}_{u}=\mathcal{A}_{u}(\mathcal{G},\mathcal{C}^{\prime}).$
The resulting model $\mathcal{G}_{u}$ should satisfy the following desiderata:
  1. (i)
@@
    Forgetting: The model should no longer be able to generate content related to targeted concepts,
    (4) $\forall c\in\mathcal{C}^{\prime},\quad\mathcal{G}_{u}(P_{text})\cap\mathcal{G}(P_{text})=\emptyset.$
  2. (ii)
    Preservation: For non-target concepts, the generative performance should be retained:
    (5) $\forall c\in\mathcal{C}\setminus\mathcal{C}^{\prime},\quad\mathcal{G}_{u}(P_{text})\approx\mathcal{G}(P_{text}).$
In practice, these constraints are relaxed using perceptual similarity metrics to quantify preservation. Specifically, the preservation condition can be approximated as:
(6) $sim\bigl{(}\mathcal{G}_{u}(P_{text}),\mathcal{G}(P_{text})\bigr{)}\geq\sigma,$
@@
where $sim(\cdot,\cdot)$ represents a perceptual similarity measure (e.g., LPIPS or CLIP score), and $\sigma$ is a predefined threshold.
3.3. Threat Model
We consider an adversary possessing white-box access to the IGM, which generates images guided by multi-modal conditioning (i.e., both text and image inputs). This IGM has been equipped with some unlearning techniques (e.g., fine-tuning (Gandikota et al., 2023; Zhang et al., 2024), closed-form modification (Gandikota et al., 2024), parameter editing (Chavhan et al., 2024), etc.) to remove some specific concepts. The adversary’s goal is to deliberately mislead the IGM to regenerate the erased content. Unlike existing attack methods restricted to textual prompts—which exhibit limited effectiveness due to the robust text-based defenses—our threat model explicitly incorporates multi-modal guidance, enabling a more comprehensive and rigorous evaluation of the unlearned model’s robustness.
3.4. Problem Formulation

### v2

@@
Abstract
Recent advances in diffusion-based image generation models (IGMs), such as Stable Diffusion (SD), have substantially improved the quality and diversity of AI-generated content. However, these models also pose ethical, legal, and societal risks, including the generation of harmful, misleading, or copyright-infringing material. Machine unlearning (MU) has emerged as a promising mitigation by selectively removing undesirable concepts from pretrained models, yet the robustness of existing methods, particularly under multi-modal adversarial inputs, remains insufficiently explored. To address this gap, we propose Recall, a multi-modal adversarial framework for systematically evaluating and compromising the robustness of unlearned IGMs. Unlike prior approaches that primarily optimize adversarial text prompts, Recall exploits the native multi-modal conditioning of diffusion models by efficiently optimizing adversarial image prompts guided by a single semantically relevant reference image. Extensive experiments across ten state-of-the-art unlearning methods and diverse representative tasks show that Recall consistently surpasses existing baselines in adversarial effectiveness, computational efficiency, and semantic fidelity to the original prompt. These results reveal critical vulnerabilities in current unlearning pipelines and underscore the need for more robust, verifiable unlearning mechanisms. More than just an attack, Recall also serves as an auditing tool for model owners and unlearning practitioners, enabling systematic robustness evaluation. Code and data are available at https://github.com/ryliu68/RECALL.
@@ Abstract
1 Introduction
@@
The emergence of image generation models (IGMs), such as Stable Diffusion (Rombach et al., 2022a), has greatly advanced the quality and diversity of AI-generated visual content. IGMs are now widely used in digital art, multimedia creation, and visual storytelling (Chen et al., 2024; Zhang et al., 2024b). However, their rapid adoption also raises serious ethical and legal concerns, particularly regarding the misuse of these models to generate harmful, misleading, or infringing content (Qu et al., 2023; Schramowski et al., 2023). Consequently, ensuring robust safety and trustworthiness mechanisms within these generative frameworks has emerged as an urgent imperative (Wang et al., 2025; 2024).
@@
Among different lines of efforts, machine unlearning (MU) has recently gained growing prominence (Zhang et al., 2024c; Park et al., 2024; Li et al., 2024b). It aims to remove sensitive concepts (e.g., nudity, violence, and copyrighted materials) from the IGMs, prohibiting the generation of sensitive or problematic content while maintaining the model’s general capability of producing benign and high-quality outputs (Schramowski et al., 2023; Kumari et al., 2023; Gandikota et al., 2024). Recent IGM unlearning (IGMU) methods utilize diverse strategies, including fine-tuning (Gandikota et al., 2023; Zhang et al., 2024a), targeted concept removal (Gandikota et al., 2024; Gong et al., 2024; Orgad et al., 2023), negative prompting (Schramowski et al., 2023), and adversarial filtering (Zhang et al., 2024c; Gong et al., 2024; Wu et al., 2025). They have proven effective in safety protection of contemporary IGMs, enforcing compliance with ethical guidelines and legal standards.
@@
Despite the rapid progress in this field, the practical robustness of these techniques is challenged, especially under adversarial scenarios. Recent studies have revealed that unlearned IGMs are still vulnerable: carefully optimized prompts can successfully circumvent safety mechanisms, prompting the unlearned models to regenerate prohibited content (Zhang et al., 2024d; Tsai et al., 2024). However, these attack methods mainly focus on perturbing the textual modality and suffer from the following critical limitations. ① Modifying textual inputs can disrupt the semantic alignment between the generated images and original prompts; ② Many approaches rely on external classifiers or additional diffusion models for adversarial text prompt optimization, incurring substantial computational overhead; ③ Their effectiveness sharply declines against robust, adversarially-enhanced unlearning methods, e.g., AdvUnlearn (Zhang et al., 2024c), RECE (Gong et al., 2024); ④ Crucially, these methods overlook the inherent multi-modal conditioning capabilities (e.g., simultaneous textual and image) of IGMs, thus missing a critical dimension of potential vulnerability.
Figure 1: Given an assumed successfully unlearned IGM $\mathcal{G}_{u}$, our adversarial image prompt $P^{adv}_{img}$ combined with the original sensitive text prompt $P_{text}$ as multi-modal guidance can circumvent the unlearning mechanism, leading to the reappearance of removed content $I^{*}$. Sensitive parts are covered by Refer to caption.
@@
Extensive empirical results conducted on ten state-of-the-art IGMU methods across four representative unlearning scenarios demonstrate that Recall consistently surpasses prior approaches in terms of adversarial effectiveness, computational efficiency, and semantic fidelity. Beyond demonstrating strong attack performance, these findings reveal critical vulnerabilities in current unlearning pipelines, underscoring their susceptibility to multi-modal guided adversarial inputs and the urgent need for more robust and verifiable unlearning mechanisms in IGMs. From the perspective of model owners, Recall can also serve as an efficient robustness auditing tool to assess the effectiveness of their unlearning procedures. Our key contributions are as follows:
@@
  • We propose Recall, the first multi-modal guided attack framework to break the robustness of IGMU techniques, allowing the protected model to regenerate unlearned sensitive concepts with high semantic fidelity.
@@
  • Recall introduces a highly efficient optimization strategy that operates solely within the unlearned model by utilizing only a single reference image, eliminating the need for auxiliary classifiers, original diffusion models, or external semantic guidance required by previous attacks.
@@
  • Through comprehensive experiments covering ten representative IGMU techniques across four diverse tasks, we empirically demonstrate the vulnerabilities of existing unlearning solutions under multi-modal attacks, revealing the urgent need for more robust safety unlearning.
2 Related Work
@@
Image Generation Models (IGMs). Diffusion-based IGMs, such as Stable Diffusion (SD) (Rombach et al., 2022b), DALL·E (OpenAI, 2023a), and Imagen (Saharia et al., 2022), have achieved impressive progress in synthesizing diverse, high-fidelity images. These models leverage large-scale datasets (e.g., LAION-5B (Schuhmann et al., 2022)) and integrate components including pre-trained text encoders (e.g., CLIP (Radford et al., 2021)), U-Net denoisers, and VAE decoders, enabling precise semantic alignment between prompts and images for a wide range of applications.
Unlearning in Image Generation Models. The proliferation of IGMs has led to increasing concerns about the generation of harmful or copyrighted content (Qu et al., 2023; Liu et al., 2025). Machine unlearning (MU) methods have been developed to selectively remove undesirable concepts from pretrained models (Kumari et al., 2023; Fan et al., 2024) while preserving overall generative capabilities. Existing IGM unlearning (IGMU) approaches can be broadly categorized as: (i) Fine-tuning-based, which update model parameters to forget specific concepts (e.g., ESD (Gandikota et al., 2023), UCE (Gandikota et al., 2024)); (ii) Guidance-based, which constrain generation at inference without modifying model weights (e.g., SLD (Schramowski et al., 2023)); and (iii) Regularization-based, which introduce forgetting objectives during training (e.g., Receler (Huang et al., 2024), FMN (Zhang et al., 2024a)). Despite their successes, these methods often exhibit limited robustness and generalization.
Adversarial Attacks on IGMU. Recent works demonstrate that adversarial prompts can circumvent IGMU defenses and recover restricted content (Li et al., 2024a; Chin et al., 2024a; Zhang et al., 2024d). White-box methods such as P4D (Chin et al., 2024a) and UnlearnDiffAtk (Zhang et al., 2024d) optimize text prompts but may suffer from high computational overhead and reduced semantic alignment, CCE (Pham et al., 2024) learns a single placeholder via textual inversion on the erased model and substitutes it at inference to recover restricted concepts, while WACE (Lu et al., 2025) regenerates forgotten content via a noise-based probe. The black-box and transfer-based approaches (Han et al., 2025; Tsai et al., 2024; Dang et al., 2025; Ma et al., 2025; Chin et al., 2024b) rely on surrogate models or textual perturbations, sometimes requiring external classifiers or access to the original diffusion model. These approaches are less effective against stronger unlearning defenses (e.g., AdvUnlearn (Zhang et al., 2024c), Receler (Huang et al., 2024)) and remain computationally intensive.
@@
Therefore, effective attack strategies should efficiently recover restricted content, preserve prompt-image semantic coherence, and exploit vulnerabilities beyond text perturbations. To this end, we propose Recall, a multi-modal adversarial framework that leverages adversarial image prompts alongside unmodified text inputs, enabling effective attacks on unlearned models via multi-modal guidance. Our method requires neither external classifiers nor access to the original IGM, making it both lightweight and effective.
3 Preliminary
3.1 Image Generation Model Unlearning
@@
Given a pretrained IGM $\mathcal{G}$ over a concept space $\mathcal{C}$, Image Generation Model Unlearning (IGMU) aims to selectively remove the model’s ability to generate content associated with a sensitive concept subset $\mathcal{C}^{\prime}\subseteq\mathcal{C}$, while preserving generative quality for the remaining concepts. Formally, an unlearning algorithm $\mathcal{A}_{u}$ produces a modified model $\mathcal{G}_{u}=\mathcal{A}_{u}(\mathcal{G},\mathcal{C}^{\prime})$. The unlearning objectives are twofold:
@@
  • Forgetting: For all $c\in\mathcal{C}^{\prime}$, the model should no longer generate content related to $c$: $\mathcal{G}_{u}(P_{text})\cap\mathcal{G}(P_{text})=\emptyset$.
@@
  • Preservation: For all $c\in\mathcal{C}\setminus\mathcal{C}^{\prime}$, the generative performance should be retained: $sim\bigl(\mathcal{G}_{u}(P_{text}),\mathcal{G}(P_{text})\bigr)\geq\sigma$, where $sim(\cdot,\cdot)$ denotes a perceptual similarity metric (e.g., CLIP score or LPIPS), and $\sigma$ is a predefined threshold.
In this work, we focus on unlearning methods and evaluation within the multi-modal, diffusion-based IGM setting, with SD as a representative backbone.
3.2 Threat Model
We consider an adversary seeking to deliberately regenerate erased content from a concept-unlearned, multi-modal (text+image) IGM. The adversary requires white-box access and the ability to invoke the model’s native multi-modal-conditioning pathway. This setting primarily targets (i) attacks: it is realistic because many applications deploy open or publicly available Stable Diffusion variants and is consistent with prior white-box threat models (Chin et al., 2024a; Pham et al., 2024; Zhang et al., 2024d). The same setup also supports (ii) unlearning red-teaming by model owners or auditors as a pre-deployment robustness assessment to locate weaknesses and guide verifiable mitigation.
3.3 Problem Formulation